Grade 10Algebra

The AP in which 4th term is -15and 9th term is -30. Find the sum of the first 10 numbers

Ankush Shailendra Giri
7 Years agoGrade 10

1 Answer

Saurabh Koranglekar
7 Years ago
Dear student

a be first term; d be common difference
Let a+3d = -15; a+8d= -30

solving these 2 eq. a= -6; d= -3

S10 = (2a+9d)*5 = (-12-27)*5 = -195